OpenStudy (anonymous):

[Particle physics] An electron and a positron undergo pair annihilation, producing two gamma ray photons: e- + e+ → 2γ If they had approximately zero kinetic energy before the annihilation, what is the wavelength of each γ ?

OpenStudy (anonymous):

Mass - Energy Equivalence

OpenStudy (anonymous):

Initial rest energy: \[E_{0} = 2(m_{0})c^2 = 1.022 \frac{ MeV }{ c^2 } = 1.637244 \times10^{-13} J\] \[E=hf=\frac{ hc }{ \lambda }\] \[\lambda = \frac{ hc }{ E } = \frac{ (4.136 \times10^{-15})(3\times10^{8}) }{ (1.637244\times10^{-13}) }\]

OpenStudy (anonymous):

I thought I was on to something but the answer was ridiculously high. I'm not sure what I've done wrong.

OpenStudy (anonymous):

your method is right.. your planks constant is wrong :P and also that energy is not of one photon, but of two photons together

OpenStudy (anonymous):

\[h = 6.63 \times 10^{-34} Js\]

OpenStudy (anonymous):

Thanks! I'll have another go.

OpenStudy (anonymous):

Awesome, I got the answer out after I divided the energy by 2. Thanks for your help!

OpenStudy (anonymous):

no problem :)